# Pragmatics in the Era of Large Language Models: A Survey on Datasets, Evaluation, Opportunities and Challenges

### v1

@@ 3 Task Types
MCQs present a scenario or question with predefined answer options, prompting selection of one or more choices. This is commonly used to evaluate LLMs by asking them to choose the correct interpretation of an utterance Hu et al. (2023); Sravanthi et al. (2024); Park et al. (2024a, b); Kim et al. (2024). It also benchmarks pretrained QA models Sap et al. (2019); Zheng et al. (2021). Answers can include scalable items or Likert scales Durmus et al. (2019); Westera et al. (2020a). MCQs simplify decision-making and improve response accuracy, making them effective for both human studies and LLM evaluations Groves (2011); Ma et al. (2024a).
Question Answering (QA).
@@
QA, including Conversational QA, replicates information-seeking dialogues with question as input and answer as output, enabling interactive communication. It evaluates reasoning about unspoken intent, a key aspect of human communication Grice (1975). QA tasks typically involve QA pairs Rashkin et al. (2019); Louis et al. (2020); Srikanth et al. (2024); Kim et al. (2024); Min et al. (2020) or multi-turn dialogues Li et al. (2023a); Zhang et al. (2018). Some datasets include both direct and literal answers to assess models’ reasoning abilities Qi et al. (2023).
Natural Language Inference (NLI).
@@
NLI usually includes a premise and a hypothesis to determine whether the hypothesis is true (entailment), false (contradiction), or undetermined (neutral) given the premise Storks et al. (2020). Pragmatic evaluations often include sentence pairs with certain different pragmatic particles or words Pedinotti et al. (2022); Koyano et al. (2022); Nizamani et al. (2024); Halat and Atlamaz (2024) or sentences from dialogues Welleck et al. (2019), testing whether models can correctly select certain pragmatic (un)related sentences. There are also datasets with NLI-like features looking at the relations of given sentences George and Mamidi (2020); Westera et al. (2020b).
Sentiment Analysis.
@@
Sentiment analysis in pragmatics, unlike analysis which focuses on polarity Gong et al. (2024), extends traditional sentiment classification by incorporating communicative intent, discourse roles, and social interaction dynamics. Datasets in this domain are designed to capture these nuances through fine-grained annotations of conversational exchanges, in topics such as emotion Buechel and Hahn (2017), sarcasm Oraby et al. (2016), hatefulness Ollagnier (2024), or medical applications for people with Autism Spectrum Disorder (ASD) Yang et al. (2021).
Image Captioning.
@@
Image captioning involves generating descriptions that convey the meaning of a given image. It is a crucial task for studying human reasoning about alternatives and efficient consideration of context Fried et al. (2023). Unlike other NLP tasks, it is inherently grounded in visual content. Datasets in this area are often designed to investigate pragmatic phenomena, such as contrastive captions and speaker-listener dynamics Tsvilodub and Franke (2023); Bao et al. (2022). Some explores negation-based descriptions to enhance pragmatic reasoning van Miltenburg et al. (2016). These datasets emphasize the interaction between visual context and textual interpretation.
Reference Games.
@@
According to Fried et al. (2023), reference games involve a speaker describing a target referent from a shared set of images, objects, or abstract illustrations to a listener, who must then identify the target. This setup is often used to model pragmatic reasoning, such as in tasks requiring context-based contrastive descriptions or multi-turn interactions Bao et al. (2022); Shaikh et al. (2023). Variants include adaptations of human games to study sequential language use Monroe et al. (2017); Khani et al. (2018) as well as visual language games in guessing given objects given a visual context de Vries et al. (2017), examining the ability to understand specific referential terms. These datasets emphasize context-dependent interpretation and collaborative communication.
Additional Tasks.
@@
There are additional task types designed specifically for certain pragmatic features, including speech-based tasks Godfrey et al. (1992); Reinig et al. (2024) for speech act studies, bias detection Kameswari et al. (2020) for assessing social norms, and various natural understanding tasks curated for a pragmatics-centered evaluation framework Sileo et al. (2022).
@@ 4 Dataset Construction
In this section, we summarize methods and data sources used in previous studies for building a pragmatics dataset. Discussing the dataset construction process is crucial for understanding the scope of current pragmatics datasets in terms of domains and source data, thus gaining insights about their limitations and special features. We categorize dataset-building approaches for evaluating pragmatic phenomena in more depth, suggesting two groups: a bottom-up approach, where source data is first collected and annotation for certain pragmatic phenomena is then performed; and a top-down approach which first determines labeling among seed data (e.g., scaler pairs), and then expands seed data to large units, such as sentences, to build up the whole dataset.
@@ 4.1 Bottom-up Approach
Source Data.
In general, there are three types of source data that current pragmatic datasets are built upon: (1) databases, such as web pages and interviews, (2) data collected directly from humans and (3) existing datasets. There is no one specific database for building a pragmatic dataset. Works focusing on dialogue or speech acts use interviews or television programs (Braga et al., 2006; George and Mamidi, 2020) as source data. Common domains for sourcing pragmatics data include open domain (Qi et al., 2023; Wang et al., 2020), debate (Durmus et al., 2019; Oraby et al., 2016), politics (Reinig et al., 2024), social media (Ollagnier, 2024; Sap et al., 2020) and law Kim et al. (2024). Fewer works directly collect data from humans due to difficulties in recruiting human participants. For instance, Yang et al. (2021) collected spoken language data from adults with high-functioning ASD to examine their pragmatic features. Other works collect data in a game setup, such as referential games, focusing on aspects such as context usage Khani et al. (2018) and pragmatic inference (Shaikh et al., 2023; Monroe et al., 2017). The majority of datasets are built upon existing datasets. They are either a unified benchmark extended on an existing pragmatic dataset (Sravanthi et al., 2024; Sileo et al., 2022; Park et al., 2024b; Guo et al., 2021), or built by annotating non-pragmatic datasets (Li et al., 2023a; Bao et al., 2022; Welleck et al., 2019; Kameswari et al., 2020; Ollagnier, 2024; van Miltenburg, 2017). Conversely, another line of research utilizes existing datasets not originally related to pragmatics to explore pragmatic phenomena. For example, Ollagnier (2024) added intent labels to analyze implicature and presupposition in hate speech.
Annotation.
Typical annotation methods involve crowd-sourced annotations (Qi et al., 2023; Li et al., 2023a; Shaikh et al., 2023) and expert annotations (Sravanthi et al., 2024; Hu et al., 2023; Park et al., 2024a). LLMs can also be used to assist annotation. For instance, to create a QA dataset featuring implicature, Srikanth et al. (2024) utilize GPT-3.5 to consolidate a list of assumptions and sub-questions annotated by experts. The model needs to either turn a sub-question into a declarative sentence to create an implicature inference, or identify the assumptions and implicature made in the sub-question. Though speeding up the annotation process and assuring data format, in the post analysis, they find using models to directly generate implicature inference is not reliable. This hybrid approach can facilitate the efficient processing of large datasets while maintaining a high level of accuracy through human participation.
@@ 4.2 Top-down Approach
Unlike the bottom-up approach where labels are created after data collection and usually provided by annotators, the labeling process in the top-down approach is motivated by linguistic theories and can be done automatically Halat and Atlamaz (2024); Nizamani et al. (2024); Koyano et al. (2022); Sap et al. (2019). For instance, to build an NLI dataset featuring implicature, Halat and Atlamaz (2024) first curated a list of scalar pairs from various linguistic categories. These scalar pairs encode logical relationships naturally, e.g., <some, all>. Then they employ GPT-4 to generate sentences based on human-annotated examples and the collected scalar pairs. Similarly, to build a pragmatic NLI dataset, Nizamani et al. (2024) collect sentences containing scalar adjectives separated by “but not”, motivated by the observation that implicatures can be explicitly reinforced Hirschberg (1985). To create premise-hypothesis pairs, they split the sentence into two parts, each containing one scalar adjective. Manual modification ensures sentence independence and grammatical correctness.
@@ 5 Evaluation
Evaluation on Model Performance.
Currently, automatic metrics dominate the evaluation landscape, with evaluation methods primarily focusing on measuring how well models align with annotated labels or golden data. While this approach is effective for tasks with well-defined categories, it often fails to capture the complexity and subtlety of pragmatic phenomena. For instance, in classification tasks, where average F1 score is widely used, many classification tasks produce binary or categorical labels that reflect only a narrow aspect of pragmatics, such as speech acts Reinig et al. (2024) or NLI Nizamani et al. (2024), without addressing broader contextual or interactive dimensions. Similarly, emotion prediction Buechel and Hahn (2017) relies on labels (e.g., “joy”, “anger”) which misses nuanced constructs such as empathy or politeness, and in some applications, the task diverges entirely from pragmatics and uses a binary label for downstream application results. Generation tasks often rely on ROUGE, BLEU, and BERTScore, as seen in cross-cultural pragmatic inference Shaikh et al. (2023) and social pragmatics Sap et al. (2020). Other metrics include factuality Qi et al. (2023), accuracy Li et al. (2023a); Kim et al. (2024), and task-specific metrics tailored to individual research goals Li et al. (2023b). While these metrics provide valuable insights into model performance, they often fall short in evaluating the broader, interconnected nature of pragmatics across tasks.
Beyond Automatic Evaluation.
Although automatic metrics like F1 score and ROUGE are useful for efficiency and standardization, they can fall short in capturing more nuanced pragmatic aspects, such as politeness. To this end, human evaluation is still needed for more comprehensive evaluation. For instance, in Yang et al. (2021), trained human annotators were asked to rate the scale of politeness and uncertainty in the spoken language data from adults with high-functioning ASD from 1 to 3. These investigated speech features cannot be easily measured by automatic evaluation metrics. Likewise, Rashkin et al. (2019) ran a crowd-sourcing human evaluation on MTurk, where annotators determined whether model-generated responses show empathy. Hence, dialogue models are evaluated not only for its general text generation ability (capturable by automatic metrics), but also more nuanced aspects in human communication, such as emotions. Such a hybrid framework provides a more complete picture of model performance, especially in terms of their capabilities of language uses beyond understanding surface meanings.
@@ 6 Opportunities and Challenges
We now discuss the opportunities and challenges identified in the surveyed resources, with a focus on the rapid development of LLMs. We focus on generalizability, the high-level evaluation outlook, alignment, as well as the potential of LLMs to advance linguistic pragmatics research both within and beyond existing resources.
@@ 6.1 Gaps and Chances in Generalizability
Current pragmatic research is dominated by English-focused datasets, limiting cross-linguistic and cultural generalizability. Out of all 57 surveyed papers, only 11 (i.e., 19%) resources include non-English languages. Though there are a few non-English resources, such as German Reinig et al. (2024) and Korean Park et al. (2024a), they focus on regional contexts, lacking critical perspectives for understanding cross-linguistic and cross-cultural pragmatics.
Data Type Diversity.
@@
Most datasets remain largely uni-modal, focusing on either text or speech, with few exploring interactions across modalities and genres. For instance, SWITCHBOARD Godfrey et al. (1992), while valuable for spoken dialogue, lacks visual data, such as gestures or facial expressions, that are essential for modeling phenomena like deixis, implicature, or politeness. A multimodal dataset combining speech, text, and video could address complex contexts like sarcasm or ambiguity resolution, aligning more closely with real-world communication and advancing pragmatic competence evaluation.
Limitations in Task Types.
@@
Pragmatic tasks typically emphasize one or a small set of phenomena, such as speech acts or emotion detection, without capturing the broader interplay between multiple pragmatic dimensions. For example, datasets like SWITCHBOARD Godfrey et al. (1992) focus on spoken dialogue and discourse markers, while the Lexical Markup Framework (LMF) Braga et al. (2006) targets lexical standardization. These tasks are often designed in isolation, making it difficult to evaluate a model’s holistic pragmatic competence. Similarly, political discourse datasets like Reinig et al. (2024) analyze specific speech acts but do not integrate additional pragmatic phenomena, such as implicature or politeness. This fragmented approach limits our ability to assess how well models handle the complex, multifaceted nature of pragmatics in real-world scenarios.
Towards Generalizability.
@@
Addressing these gaps requires the creation of multilingual, multimodal datasets that incorporate visual, textual, and spoken interactions. Leveraging hybrid approaches that combine synthetic LLM-generated data with human validation could provide scalable and diverse benchmarks. In addition, pragmatic task designs should move beyond uni-modal classification and generation to explore holistic, real-world applications that evaluate models’ ability to integrate multiple pragmatic phenomena, such as emotion detection, politeness, and implicature resolution, within the same task framework. These advancements will enable a more comprehensive and inclusive understanding of pragmatic competence in computational models.

### v3

@@ 3 Task Types
Multiple-Choice Question (MCQ) Setup. MCQs present a scenario or question with predefined answer options, prompting selection of one or more choices. This is commonly used to evaluate LLMs by asking them to choose the correct interpretation of an utterance Hu et al. (2023); Sravanthi et al. (2024); Park et al. (2024a, b); Kim et al. (2024). It also benchmarks pretrained QA models Sap et al. (2019); Zheng et al. (2021). Answers can include scalable items or Likert scales Durmus et al. (2019). MCQs simplify decision-making and improve response accuracy, making them effective for both human studies and LLM evaluations Groves (2011); Ma et al. (2024a).
@@
Question Answering (QA). QA, including Conversational QA, models information-seeking dialogues by mapping questions to answers, enabling interactive communication. It evaluates reasoning about unspoken intent, a key aspect of human communication Grice (1975). QA tasks typically involve QA pairs Rashkin et al. (2019); Louis et al. (2020); Srikanth et al. (2024); Kim et al. (2024); Min et al. (2020) or multi-turn dialogues Li et al. (2023a); Zhang et al. (2018). Some datasets include direct or literal answers to assess models’ reasoning abilities Qi et al. (2023); Miao et al. (2024).
@@
Natural Language Inference (NLI). NLI usually includes a premise and a hypothesis to determine whether the hypothesis is true (entailment), false (contradiction), or undetermined (neutral) given the premise Storks et al. (2020). Pragmatic evaluations often include sentence pairs with certain different pragmatic particles or words Jeretic et al. (2020); Pedinotti et al. (2022); Koyano et al. (2022); Nizamani et al. (2024); Halat and Atlamaz (2024) or sentences from dialogues Welleck et al. (2019), testing whether models can correctly select certain pragmatic (un)related sentences. There are also datasets with NLI-like features looking at the relations of given sentences George and Mamidi (2020); Westera et al. (2020); Cong (2024).
@@
Sentiment Analysis. Sentiment analysis in pragmatics, unlike analysis which focuses on polarity Gong et al. (2024), extends traditional sentiment classification by incorporating communicative intent, discourse roles, and social interaction dynamics. Datasets in this domain are designed to capture these nuances through fine-grained annotations of conversational exchanges, in topics such as emotion Buechel and Hahn (2017), sarcasm Oraby et al. (2016), hatefulness Ollagnier (2024), or medical applications for people with Autism Spectrum Disorder (ASD, Yang et al. 2021).
@@
Image Captioning. Image captioning involves generating descriptions that convey the meaning of a given image. It is a crucial task for studying human reasoning about alternatives and efficient consideration of context Fried et al. (2023). Unlike other NLP tasks, it is inherently grounded in visual content. Datasets in this area are often designed to investigate pragmatic phenomena, such as contrastive captions and speaker-listener dynamics Tsvilodub and Franke (2023); Bao et al. (2022). Some explores negation-based descriptions to enhance pragmatic reasoning van Miltenburg et al. (2016). These datasets emphasize the interaction between visual context and textual interpretation.
@@
Reference Games. According to Fried et al. (2023), reference games involve a speaker describing a target referent from a shared set of images, objects, or abstract illustrations to a listener, who must then identify the target. This setup is often used to model pragmatic reasoning, such as in tasks requiring context-based contrastive descriptions or multi-turn interactions Bao et al. (2022); Shaikh et al. (2023). Variants include adaptations of human games to study sequential language use Monroe et al. (2017); Khani et al. (2018) as well as visual language games in guessing given objects given a visual context De Vries et al. (2017); Takmaz et al. (2020, 2023); Greco et al. (2023), examining the ability to understand specific referential terms. These datasets emphasize context-dependent interpretation and collaborative communication.
@@
Additional Tasks. There are additional task types designed specifically for certain pragmatic features, including speech-based tasks Godfrey et al. (1992); Reinig et al. (2024) for speech act studies, bias detection Kameswari et al. (2020) for assessing social norms, cloze-style tasks for discourse markers Pandia et al. (2021); Sadlier-Brown et al. (2024), and various natural understanding tasks curated for a pragmatics-centered evaluation framework Sileo et al. (2022).
@@ 4 Dataset Construction
In this section, we summarize methods and data sources used in previous studies for building a pragmatic dataset. Discussing the dataset construction process is crucial for understanding the scope of current pragmatic datasets in terms of domains and source data, thereby gaining insights about their limitations and special features. We categorize dataset-building approaches for evaluating pragmatic phenomena in more depth, suggesting two groups: a bottom-up approach, where source data is first collected and annotation for certain pragmatic phenomena is then performed; and a top-down approach which first determines labeling among seed data (e.g., scalar pairs), and then expands seed data to large units, such as sentences, to build up the whole dataset.
@@ 4.1 Bottom-up Approach
Source Data. In general, there are three types of source data that current pragmatic datasets are built upon: (1) databases, such as web pages and interviews, (2) data collected directly from humans and (3) existing datasets.
There is no one specific database for building a pragmatic dataset. Works focusing on dialogue or speech acts use interviews or television programs (Braga et al., 2006; George and Mamidi, 2020) as source data. Common domains for sourcing pragmatic data include open domain (Qi et al., 2023; Wang et al., 2020), debate (Durmus et al., 2019; Oraby et al., 2016), politics (Reinig et al., 2024), social media (Ollagnier, 2024; Sap et al., 2020) and law Kim et al. (2024). Fewer works collect data from humans due to difficulties in recruiting human participants. For instance, Yang et al. (2021) collected spoken language data from adults with high-functioning ASD to examine their pragmatic features. Other works collect data in a game setup, such as reference games, focusing on aspects such as context usage Khani et al. (2018) and pragmatic inference (Shaikh et al., 2023; Monroe et al., 2017). The majority of datasets are built upon existing datasets. They are either a unified benchmark extended on an existing pragmatic dataset (Sravanthi et al., 2024; Sileo et al., 2022; Park et al., 2024b; Guo et al., 2021), or built by annotating non-pragmatic datasets (Li et al., 2023a; Bao et al., 2022; Welleck et al., 2019; Kameswari et al., 2020; Ollagnier, 2024; van Miltenburg, 2017). Conversely, another line of research uses existing datasets not originally related to pragmatics to explore pragmatic phenomena. For example, Ollagnier (2024) adds intent labels to analyze implicature in hate speech.
Annotation. Typical annotation methods involve crowd-sourced annotations (Qi et al., 2023; Li et al., 2023a; Shaikh et al., 2023) and expert annotations (Jeretic et al., 2020; Sravanthi et al., 2024; Hu et al., 2023; Park et al., 2024a). LLMs can also be used to assist annotation. For instance, to create a QA dataset featuring implicature, Srikanth et al. (2024) utilize GPT-3.5 to consolidate a list of assumptions and sub-questions annotated by experts. The model needs to either turn a sub-question into a declarative sentence to create an implicature inference, or identify the assumptions and implicature made in the sub-question. Though speeding up the annotation process and assuring data format, in the post analysis, they find using LLMs to directly generate implicature inference is not reliable. Besides, Cong (2024) employs ChatGPT in generating synthetic stimuli for implicature sentences, with human raters reviewing after the automatic generation process, with the setup of Likert scaling scoring. These hybrid approaches can facilitate the efficient processing of large datasets while maintaining a high level of accuracy through human participation.
@@ 4.2 Top-down Approach
Unlike the bottom-up approach, where labels are created after data collection and usually provided by annotators, the labeling process in the top-down approach is motivated by linguistic theories and can be done automatically Halat and Atlamaz (2024); Nizamani et al. (2024); Koyano et al. (2022); Sap et al. (2019). For instance, to build an NLI dataset featuring implicature, Halat and Atlamaz (2024) first curate a list of scalar pairs from various linguistic categories. These scalar pairs encode logical relationships naturally, e.g., <some, all>. Then they employ GPT-4 to generate sentences based on human-created examples as demonstrations (I read some of the books. v.s. I read all of the books.). Similarly, to build a pragmatic NLI dataset, Nizamani et al. (2024) collect sentences containing scalar adjectives separated by “but not” (e.g., good but not great), motivated by the observation that implicatures can be explicitly reinforced Hirschberg (1985). To create premise-hypothesis pairs, they split the sentence into two parts, each containing one scalar adjective. For instance, the original sentence: “These shows were good, but not great.” can be split into: “These shows were good.” and “These shows were great.” Manual modification ensures sentence independence and grammatical correctness.
@@ 5 Evaluation
Evaluation on Model Performance. Currently, automatic metrics dominate the evaluation landscape, with evaluation methods primarily focusing on measuring how well models align with annotated labels or golden data. While this approach is effective for tasks with well-defined categories, it often fails to capture the complexity and subtlety of pragmatic phenomena. For instance, in classification tasks, where average F1 score is widely used, many classification tasks produce binary or categorical labels that reflect only a narrow aspect of pragmatics, such as speech acts Reinig et al. (2024) or NLI Nizamani et al. (2024), without addressing broader contextual or interactive dimensions. Similarly, emotion prediction Buechel and Hahn (2017) relies on labels (e.g., “joy”, “anger”) which misses nuanced constructs such as empathy or politeness, and in some applications, the task diverges entirely from pragmatics and uses a binary label for downstream application results. Generation tasks often rely on ROUGE, BLEU, and BERTScore, as seen in cross-cultural pragmatic inference Shaikh et al. (2023) and social pragmatics Sap et al. (2020). Other metrics include factuality Qi et al. (2023), accuracy Li et al. (2023a); Kim et al. (2024), and task-specific metrics tailored to individual research goals Li et al. (2023b). While these metrics provide valuable insights into model performance, they often fall short in evaluating the broader, interconnected nature of pragmatics across tasks.
Beyond Automatic Evaluation. Although automatic metrics like F1 score and ROUGE are useful for efficiency and standardization, they can fall short in capturing more nuanced pragmatic aspects, such as politeness. To this end, human evaluation is still needed for more comprehensive evaluation. For instance, in Yang et al. (2021), trained human annotators were asked to rate the scale of politeness and uncertainty in the spoken language data from adults with high-functioning ASD from $11$ to $33$. These investigated speech features cannot be easily measured by automatic evaluation metrics. Likewise, Rashkin et al. (2019) ran a crowd-sourcing human evaluation on MTurk, where annotators determined whether model-generated responses show empathy. Hence, dialogue models are evaluated not only for its general text generation ability (capturable by automatic metrics), but also more nuanced aspects in human communication, such as emotions. Such a hybrid framework provides a more complete picture of model performance, especially in terms of their capabilities of language uses beyond understanding surface meanings.
@@ 6 Opportunities and Challenges
We now discuss the opportunities and challenges identified in the surveyed resources, with a focus on the rapid development of LLMs. We focus on generalizability, the high-level evaluation outlook, alignment, as well as the potential of LLMs to advance research in linguistic pragmatics both within and beyond existing resources.
@@ 6.1 Gaps and Chances in Generalizability
English-Centric Bias. Current pragmatic research is dominated by English-focused datasets, limiting cross-linguistic and cultural generalizability. Out of all $5757$ surveyed papers, only $1111$ (i.e., 19191919%) resources include non-English languages. Though there are a few non-English resources, such as German Reinig et al. (2024) and Korean Park et al. (2024a), they focus on regional contexts, lacking critical perspectives for understanding cross-linguistic and cross-cultural pragmatics.
Human/Demographic Bias. NLP frameworks that acknowledge that in certain tasks human labels and judgments may be inherently subjective, and display a lot of variation between subjects (Plank, 2022; Cabitza et al., 2023; Kern et al., 2023, see e.g.,). In addition, recently, more and more works have focused on simulating human behaviours and opinions including social demographic profiles of real human participants (Argyle et al., 2023, e.g.,). The demographics aspect is in general an important factor in the data collection phase in terms of human label variation, and should be considered while conducting LLM evaluations, which we see as a gap in current works.
@@
Data Type Diversity. Most datasets remain largely uni-modal, focusing on either text or speech, with few exploring interactions across modalities and genres. For instance, SWITCHBOARD Godfrey et al. (1992), while valuable for spoken dialogue, lacks visual data, such as gestures or facial expressions, that are essential for modeling phenomena like deixis, implicature, or politeness. A multimodal dataset combining speech, text, and video could address complex contexts like sarcasm or ambiguity resolution, aligning more closely with real-world communication and advancing pragmatic competence evaluation.
@@
Limitations in Task Types. Pragmatic tasks typically emphasize one or a small set of phenomena, such as speech acts or emotion detection, without capturing the broader interplay between multiple pragmatic dimensions. For example, datasets like SWITCHBOARD Godfrey et al. (1992) focus on spoken dialogue and discourse markers, while the Lexical Markup Framework Braga et al. (2006) targets lexical standardization. These tasks are often designed in isolation, making it difficult to evaluate a model’s holistic pragmatic competence. Similarly, political discourse datasets like Reinig et al. (2024) analyze specific speech acts but do not integrate additional pragmatic phenomena, such as implicature or politeness. This fragmented approach limits our ability to assess how well models handle the complex, multifaceted nature of pragmatics in real-world scenarios.
@@
Towards Generalizability. Addressing these gaps requires the creation of multilingual, multimodal datasets that incorporate visual, textual, and spoken interactions. Leveraging hybrid approaches that combine synthetic LLM-generated data with human validation could provide scalable and diverse benchmarks. In addition, pragmatic task designs should move beyond uni-modal classification and generation to explore holistic, real-world applications that evaluate models’ ability to integrate multiple pragmatic phenomena, such as emotion detection, politeness, and implicature resolution, within the same task framework. These advancements will enable a more comprehensive and inclusive understanding of pragmatic competence in computational models.